OpenStudy (anonymous):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 |5x + 6| = 41 Equation 2 |2x + 13| = 28

OpenStudy (anonymous):

I get X=7 and X=-47/5 for the first one and X=15/2 and X=-41/2. No common solution.